### v1

@@ Abstract
First, we theoretically model LST-liquidity on AMMs. This includes categorizing suitable AMM types for LST liquidity, as well as deriving formulas for the necessary returns from trading fees to adequately compensate liquidity providers under the particular price trajectories of LSTs. Two relevant metrics are considered. Firstly, losses compared to holding the liquidity outside the AMM (loss-versus-holding, or “impermanent loss”). Secondly, the relative profitability compared to fully staking the capital (loss-versus-staking) which is a metric is specifically introduced for the case of LST-liquidity.
Subsequently, we empirically measure these metrics for Ethereum LSTs across the most relevant AMM pools. We find that, while trading fees often compensate for impermanent loss, fully staking is more profitable for many pools, putting the sustainability of current LST allocation to AMMs into question.
@@ 1 Introduction
In traditional staking on proof-of-stake (PoS) blockchains, stakers lock (“stake”) native blockchain tokens with a validator, typically subject to a lock-up period. Validators maintain the blockchain network and receives staking rewards in exchange. Stakers have the option to run their own validators or become customers of a staking pool operator. In the case of the Ethereum blockchain, the minimum collateral of 32 Ether (ETH) is required to set up and operate a validator.
In contrast, liquid staking tokens (LSTs) allow for more simple and flexible staking. LSTs are tokenized representations of the staked underlying asset, and can be minted and redeemed for the corresponding amount of the underlying asset directly from the LST protocol. Additionally, since LSTs are fungible tokens, they can be bought and sold on exchanges, in particular on decentralized automated market makers (AMMs).
@@
The emergence of LSTs has already transformed decentralized lending, with LSTs becoming the dominant collateral [9], and has also reshaped investment strategies within DeFi. A unique advantage and feature of LSTs which is also vital to the intended functioning of LSTs is the possibility to trade them on AMMs, which enables instant entering or exiting of staking positions. Precisely this aspect, i.e. LST liquidity on AMMs, is the subject of this research.
@@ 1.1 Methodology and Contribution
To the best of our knowledge, this is the first work to systematically study – both theoretically and empirically – the functioning of LST-liquidity on AMMs. A central aspect in this regard is the profitability of providing LSTs as liquidity to AMMs.
On AMMs, LSTs are commonly traded against their underlying asset. Since the price of LSTs is by design either pegged to a unit of the underlying asset or experiences a constant increase compared to the underlying asset as a result of staking rewards, such LST trading pairs exhibit a specific behavior: Apart from short-term price fluctuations, their price changes are constant and predictable.
This paper studies AMM liquidity in this particular setting. We begin by categorizing suitable AMM types for LST liquidity. Subsequently, we concentrate on LP profitability. We analyze empirically whether, historically, trading fees sufficiently compensated liquidity provider (LPs) for the costs of providing liquidity. Moreover, we theoretically derive formulas for the necessary returns from trading fees, as a function of the staking rate.
For this purpose, we consider the following two metrics measuring the losses to LPs on AMMs.
@@
A second relevant measure, that we introduce specifically for LSTs, is comparing proving LST liquidity on an AMM to staking. We coin the term loss-versus-staking (LVS) for this metric. Instead of providing an LST and its underlying assets as liquidity on an AMM, LP can stake its full capital. In the latter case, the LP earns staking rewards on their full capital compared to trading fees from the AMM and staking rewards on part of their capital in the first case.
This work begins theoretically by deriving formulas for the necessary returns from trading fees required to adequately compensate LPs for each of the losses under the particular price trajectory of LSTs. This results in expressions for the required returns as a function of the staking rate $r$.
Subsequently, we empirically measure these metrics for the largest Ethereum LSTs: stETH and wstETH by Lido [2], Coinbase’s cbETH [3] and Rocketpool’s rETH [11]. Together, they represent over 90% of TVL in liquid staking [^1] https://defillama.com/lsd. For these tokens, we consider the largest AMM pools in terms of TVL, which are on the decentralized exchanges Curve (v1 [22] and v2 [23]) and Uniswap (v3 [17]) on the Ethereum blockchain [21].
We find that simply holding LSTs is more profitable than allocating LSTs to AMM liquidity pools: trading fees on AMMs did not compensate LPs for not staking their capital fully (by allocating it to LSTs) during the studied time period. Yet, impermanent loss was sufficiently compensated for most of the analyzed liquidity pools. Nonetheless, our result put the sustainability of the current allocation of LST liquidity on AMMs into question.
@@ 2 Related Work
On the empirical side, the general profitability of liquidity providers on AMMs has been studied in the literature for different AMMs, token pairs, and time periods: for Uniswap v2 in [28] and for Uniswap v3 in [24, 27, 29]. Most of these works focus on impermanent loss (loss-versus-holding) as a measure of LP profitability. As an alternative metric, loss-versus-rebalancing (LVR) that benchmarks AMM liquidity against a rebalancing portfolio that hedges the position’s market risk was introduced in [31], [30].
@@ 3.1 Liquid Staking Tokens
LSTs are tokenized representations of staked native PoS blockchain tokens. They can be redeemed for the underlying token at any time (possibly with a delay) at the LST protocol, or traded on AMMs. Consequently, their market price on AMMs oscillates around the value of the staked assets they represent. Like other pegged tokens such as stablecoins (pegged to fiat currency) and wrapped tokens (pegged to tokens from other blockchains), LSTs offer a means to represent staked tokens in a liquid and tradable form. LSTs can be categorized into two main types based on their staking distribution models: rebase or reward [25]. The third model - dual token LSTs - is disappearing [12]. LSTs also vary in validator selection processes and governance [26]. Additionally, liquid staking protocols typically charge a fee, usually around 10-25% of the staking rewards, to cover operational costs and provide incentives for the platform’s maintenance and development.
3.1.1 Rebase-LST
@@ 3.3 Constant Function Market Makers (CFMMs)
Constant Product (CPMM)
@@
Constant Product with Concentrated Liquidity (CLMM)
@@ 4 The Theory of LSTs on AMMs
While LSTs can potentially be paired and traded against any other token on AMMs, the majority of LST liquidity lies in liquidity pools containing the LST and the underlying token. This supports the main purpose of LSTs on AMMs: Entering or exiting a staking positions by buying or selling the LST for the underlying. Pairing LSTs against their underlying is also advantageous, since the volatility between the two should be low (if the LST functions well).
@@ 4.1 Suitable types of AMMs for LSTs
Different types of AMMs come with different properties and limitations, rendering them more or less well suited for specific liquid staking token pairs. Refer to Table 1 for which AMMs are most suitable for rebase and reward LSTs to achieve high capital efficiency for LPs and low price impact for traders. While LSTs could potentially be traded on CPMMs, their low capital inefficiency for assets that move little in price, such as LSTs, makes them an unattractive choice.
Rebase-LSTs maintain a price pegged to 1 ETH, aligning perfectly with the Stableswap invariant, which was specifically designed for pairs with stable prices. Although CLMMs are theoretically also well-suited for trading pairs with stable prices, they are currently not widely used for trading rebase-LSTs due to the technical incompatibility of Uniswap v3 – the market-leading CLMM - with rebase-LSTs [7].
@@
rebase-LST
ETH
Stableswap
rebase-LST
rebase-LST
Stableswap
reward-LST
ETH
Cryptoswap, CLMM(*)
reward-LST
reward-LST
Cryptoswap, CLMM
@@ 4.2 Losses and Required Returns for LPs
4.3 LST prices
Reward LSTs
@@ 4.3 LST prices
In later sections (starting Section 4.4), we derive expressions for LP losses under this price process for certain AMMs.
Rebase LSTs
For trading pairs between rebase LSTs and the underlying asset, the matter is more straightforward, as the price of rebase LSTs is pegged to the price of the underlying asset. We only briefly discuss this case at a high level.
Finally, the value of the fully staked portfolio is $V_{LST}(t)=V_{LP}(0)e^{rt}$, leading to
$rr_{LVS}(t)=\dfrac{V_{LP}(0)e^{rt}}{V_{LP}(0)(1/2+e^{rt}/2)}-1=\dfrac{2e^{rt}}{1+e^{rt}}.$
4.4 Required returns for CPMM
@@
For the ideal LST price trajectory with $\sigma=0$, this implies $\mathbb{E}[rr_{LVH}]=(e^{r/2}+e^{-r/2})/2-1$.
For loss-versus-staking, we derive the following expression in Appendix 0.B:
@@
4.5 Required Returns for Concentrated Liquidity CPMMs
@@ 5.2 Analysis
We begin the analysis with the stETH-ETH pool on Curve v1 - the largest pool in terms of TVL. The development of the considered portfolio values – LP (including fees), HOLD, and LST (full LST position) – as well as the reference staking rate are depicted in Figure 2. We additionally plot the wealth of an LP including CRV rewards (LP+CRV). In Curve, LPs earn additional rewards for liquidity provision on top of trading fees. These rewards are distributed in Curve governance tokens - CRV according to its emissions [14]. While we observe that the additional CRV rewards did not significantly impact the wealth of LPs in this pool, they do make a difference in other pools (see Figure 8 in the appendix).
@@
We find that during most of the analyzed periods, liquidity provision resulted in both a loss-versus-holding and a loss-versus-staking over a 7-day period during about half the observation period. Figure10 in the appendix, displaying the same comparison, but for a 30-day moving average, shows a similar picture.
Impact of CRV Rewards
@@
LST pools on Curve
Figures 8, 9 and 10 show the development of the wealth of LP, HOLD, and LST, as well as the 7-day and 30-day moving averages of their differences, for the five LST pools on Curve.
Three of the pools allow swapping stETH for ETH but differ in the parameters used and, consequently, in the behavior of the bonding curve. While the leverage parameter A for the largest pool in terms of TVL is 30, the “concentrated” (A=1000) and “ng” (A=1500) pools that have later inception dates, offer more concentrated liquidity. Figure 10 shows that the more concentrated pools almost always compensate LPs for loss-versus-holding, but trading fees are not sufficient to compensate for loss-versus-staking.
The analysis also includes two pools with reward-LSTs – rETH and cbETH – deployed to Curve v2. As reward-LSTs, rETH and cbETH are not pegged to 1 ETH and their pools use the Cryptoswap Invariant - an AMM that concentrates liquidity around their current prices. The returns in the rETH and cbETH pools were significantly impacted by the distribution of CRV token rewards. As depicted in Figure 8 (blue dotted lines), the LP wealth including CRV tokens actually significantly exceeded the staking returns for rETH and cbETH. Without rewards, the picture is similar
LST pools on Uniswap
Figures 11, 12 and 13 show the wealth of LP, HOLD, and LST, as well as 7-day and 30-day moving average of wealth difference, for the four LST pools on Uniswap v3. Due to incompatibility with rebase tokens, all LSTs listed at Uniswap are reward-based (wstETH, rETH, and cbETH). The moving average for the weekly positions in Figure 12 shows that the pools with the higher fee rate of 5 bp were able to generate sufficient amount of fees to compensate for LVS during certain time periods, while the 1 bp wstETH-ETH pool consistently failed to achieve this.
Over the whole observation period, total fee returns were generally high enough to compensate LPs for LVS in the 5bp pools, especially in the more recent past (see right plots in Figure 11). However, the actual profitability of liquidity positions depends on their holding period, as it is affected by price fluctuations of cbETH and rETH. In particular, depositing into the pool during the depegs following the FTX insolvency in November 2022 (rETH being overpriced while cbETH was underpriced), can lead to different results.
@@ 5.3 Discussion
CRV rewards
We find that most Curve pools currently are only profitable for LPs, especially compared to fully staking, when taking the CRV token rewards to LPs into account. The amounts of CRV token rewards emitted to pools are decided by the Curve governance, and can change every week. A drop in their distribution to the LST pools, could lead to lower returns for LPs and subsequently a decrease of liquidity in these pools.
CLMM Rebalancing
@@
Reward-LST token pairs
@@
6 Conclusions
Liquid staking tokens as liquidity on AMMs has grown to a vital part of the DeFi landscape, but has so far seen liquidated academic attention. This work systemizes liquidity provision of LSTs on AMMs, and both theoretically and empirically studies its returns.
In addition to LVH and LVR, the two most commonly-used metrics to measure LP losses on AMM, we introduce loss-versus-staking (LVS) which is specifically tailored to the case of LST pools on AMMs.
Our empirical results show that most AMM pools with LSTs have historically not compensated for loos-versus-staking. Nevertheless, most pools did compensate for LVH (“impermanent loss”).

### v2

@@ Abstract
First, we model LST-liquidity on AMMs theoretically, categorizing suitable AMM types for LST liquidity and deriving formulas for the necessary returns from trading fees to adequately compensate liquidity providers under the particular price trajectories of LSTs. For the latter, two relevant metrics are considered: (1) losses compared to holding the liquidity outside the AMM (loss-versus-holding, or “impermanent loss”), and (2) the relative profitability compared to fully staking the capital (loss-versus-staking) which is specifically tailored to the case of LST-liquidity.
Next, we empirically measure these metrics for Ethereum LSTs across the most relevant AMM pools. We find that, while trading fees often compensate for impermanent loss, fully staking is more profitable for many pools, raising questions about the sustainability of the current LST liquidity allocation to AMMs.
@@ 1 Introduction
In traditional staking on proof-of-stake (PoS) blockchains [29], stakers lock (“stake”) native blockchain tokens with a validator, typically subject to a lock-up period. Validators maintain the blockchain network and receive staking rewards in exchange. Stakers have the option to run their own validators or become customers of a staking pool operator. In the case of the Ethereum blockchain, the minimum collateral of 32 Ether (ETH) is required to set up and operate a validator.
In contrast, liquid staking tokens (LSTs) allow for more simple and flexible staking. LSTs are tokenized representations of the staked underlying asset, and can be minted and redeemed for the corresponding amount of the underlying asset directly from the LST protocol. [^1] Some blockchains, such as Cardano, come with “native liquid staking” without the need for LST protocols. Additionally, since LSTs are fungible tokens, they can be bought and sold on exchanges, in particular on decentralized exchanges running automated market maker protocols (AMMs).
@@
In particular, the emergence of LSTs has transformed decentralized lending, with LSTs becoming the dominant collateral [21], and has also reshaped investment strategies within DeFi. A unique advantage and feature of LSTs which is vital to their intended functionality is the possibility to trade them on AMMs, which enables instant entering or exiting of staking positions. Precisely this aspect, i.e. LST liquidity on AMMs, is the subject of this research.
We systematically study – both theoretically and empirically – the functioning of LST-liquidity on AMMs. A central aspect in this regard is the profitability of providing LSTs as liquidity to AMMs.
On AMMs, LSTs are commonly traded against their underlying asset. Since the price of LSTs is by design either pegged to a unit of the underlying asset or experiences a constant increase compared to the underlying asset as a result of staking rewards, such LST trading pairs exhibit a specific behavior: Apart from short-term price fluctuations, their price changes are constant and predictable. This paper studies AMM liquidity in this particular setting.
@@ 1.1 Methodology and Contribution
We begin by categorizing suitable AMM types for LST liquidity. Subsequently, we concentrate on the profitability of liquidity providers (LPs). We theoretically derive formulas for the necessary returns from trading fees as a function of the staking rate. For this purpose, we consider the following two metrics measuring the losses to LPs on AMMs.
@@
A second relevant measure, that we introduce specifically for LSTs, is comparing providing LST liquidity on an AMM to fully staking the capital. We coin the term loss-versus-staking (LVS) for this metric. Instead of providing an LST and its underlying assets as liquidity on an AMM, LPs can stake their capital by fully holding it in LSTs. In the latter case, LPs earn staking rewards on their full capital compared to trading fees from the AMM and staking rewards on part of their capital in the first case.
For each metric, we derive the formula for the necessary returns from trading fees required to adequately compensate LPs for each of the losses under the particular price trajectory of LSTs. This results in expressions for the required returns as a function of the staking rate $r$.
Subsequently, we empirically measure these metrics for the largest Ethereum LSTs: stETH and wstETH by Lido [23], Coinbase’s cbETH [9] and Rocketpool’s rETH [28]. Together, they represent over 80% of TVL in liquid staking [^2] https://defillama.com/lsd. For these tokens, we consider the largest AMM pools in terms of TVL, which are on the decentralized exchanges Curve (v1 [11] and v2 [12]) and Uniswap (v3 [4]) on the Ethereum blockchain [8].
We find that simply holding LSTs is more profitable than allocating LSTs to most AMM liquidity pools: trading fees on the AMMs did not compensate LPs for not staking their capital fully (by allocating it to LSTs) during the studied time period. Yet, impermanent loss was sufficiently compensated for most of the analyzed liquidity pools. Nonetheless, our result put the sustainability of the current allocation of LST liquidity on AMMs into question. This is particularly noteworthy since the amount of AMM liquidity is already relatively low compared to the overall supply of LSTs at the time of writing (\$0.5 billion liquidity compared to more than \$30 billion supply, cf. Figure 7).
@@ 2 Related Work
On the empirical side, the general profitability of liquidity providers on AMMs has been studied in the literature for different AMMs, token pairs, and time periods: for Uniswap v2 in [19] and for Uniswap v3 in [14, 18, 25]. Most of these works focus on impermanent loss (loss-versus-holding) as a measure of LP profitability. As an alternative metric, loss-versus-rebalancing (LVR) that benchmarks AMM liquidity against a rebalancing portfolio that hedges the position’s market risk was introduced in [27, 26], and used to measure LP profitability empirically in [15].
@@ 3.1 Liquid Staking Tokens
LSTs are tokenized representations of staked native PoS blockchain tokens. They can be redeemed for the underlying token at any time (possibly with a delay) at the LST protocol, or traded on AMMs. Consequently, their market price on AMMs oscillates around the value of the staked assets they represent. Like other pegged tokens such as stablecoins (most commonly pegged to fiat currency) and wrapped tokens (pegged to tokens from other blockchains), LSTs offer a means to represent staked tokens in a liquid and tradable form. LSTs can be categorized into two main types based on their staking distribution models: rebase or reward [16]. The third model - dual token LSTs - is disappearing [31]. LSTs also vary in validator selection processes and governance [17]. Additionally, liquid staking protocols typically charge a fee, usually around 10-25% of the staking rewards, to cover operational costs and provide incentives for the platform’s maintenance and development.
3.1.1 Rebase-LST.
@@ 3.3 Constant Function Market Makers (CFMMs)
3.3.1 Constant Product (CPMM).
@@
3.3.2 Constant Product with Concentrated Liquidity (CLMM).
@@ 4 The Theory of LSTs on AMMs
While LSTs can potentially be paired and traded against any other token on AMMs, the majority of LST liquidity lies in liquidity pools containing the LST and the underlying token. This supports the main purpose of LSTs on AMMs: Entering or exiting staking positions by buying or selling the LST for the underlying. Pairing LSTs against their underlying is also advantageous, since the volatility between the two should be low (if the LST functions well).
@@ 4.1 Suitable types of AMMs for LSTs
Different types of AMMs come with different properties and limitations, rendering them more or less well suited for specific liquid staking token pairs. Refer to Table 1 for which AMMs are most suitable for rebase and reward LSTs to achieve high capital efficiency for LPs and low price impact for traders. While LSTs could potentially be traded on CPMMs, their low capital efficiency for assets that move little in price, such as LSTs, makes them an unattractive choice.
Rebase-LSTs maintain a price pegged to 1 ETH, aligning perfectly with the Stableswap invariant, which was specifically designed for pairs with stable prices. Although CLMMs are theoretically also well-suited for trading pairs with stable prices, they are currently not widely used for trading rebase-LSTs due to the technical incompatibility of Uniswap v3 – the market-leading CLMM - with rebase-LSTs [35, 24].
@@
rebase-LST ETH Stableswap
rebase-LST rebase-LST Stableswap
reward-LST ETH Cryptoswap, CLMM(*)
reward-LST reward-LST Cryptoswap, CLMM
@@ 4.2 Losses and Required Returns for LPs
4.3 Required returns for rebase LSTs
For trading pairs between rebase LSTs and the underlying asset, the matter is relatively straightforward, as the price of rebase LSTs is pegged to the price of the underlying asset. We only briefly discuss this case at a high level.
Finally, the value of the fully staked portfolio is $V_{LST}(t)=V_{LP}(0)e^{rt}$, implying
$rr_{LVS}(t)=\dfrac{V_{LP}(0)e^{rt}}{V_{LP}(0)(1/2+e^{rt}/2)}-1=\dfrac{2e^{rt}}{1+e^{rt}}-1.$
4.4 Required returns for reward LSTs
@@
In the following, we derive expressions for LP losses under this price process for certain AMMs.
4.4.1 Required returns for CPMMs.
@@
For the ideal LST price trajectory with $\sigma=0$, this implies $\mathbb{E}[rr_{LVH}]=(e^{r/2}+e^{-r/2})/2-1$. For loss-versus-staking, we derive the following expression in Appendix 0.B:
@@
4.4.2 Required Returns for Concentrated Liquidity CPMMs.
@@ 5.2 Analysis
We begin the analysis with the stETH-ETH pool on Curve v1 - the largest pool in terms of TVL. The development of the considered portfolio values – LP (including fees), HOLD, and LST (full LST position) – as well as the reference staking rate are depicted in Figure 2. We additionally plot the wealth of an LP including CRV rewards (LP+CRV). In Curve, LPs earn additional rewards for liquidity provision on top of trading fees. These rewards are distributed in Curve governance tokens - CRV according to its emissions [10]. While we observe that the additional CRV rewards did not significantly impact the wealth of LPs in this pool, they do make a difference in other pools (see Figure 8(a) in the appendix).
@@
We find that during most of the analyzed periods, liquidity provision resulted in both a loss-versus-holding and a loss-versus-staking over a 7-day period during about half the observation period. Figure 9(a) in the appendix, displaying the same comparison, but for a 30-day moving average, shows a similar picture.
5.2.2 Impact of CRV Rewards.
@@
5.2.3 LST pools on Curve.
Figures8(a) and 9(a) exhibit the wealth of LP, HOLD, and LST, as well as the 30-day moving averages of their differences, for the five LST pools on Curve. Three of the pools allow swapping stETH for ETH but differ in the parameters used and, consequently, in the behavior of the bonding curve. While the leverage parameter A for the largest pool in terms of TVL is 30, the “concentrated” (A=1000) and “ng” (A=1500) pools that have later inception dates, offer more concentrated liquidity. Figure 9(a) shows that the more concentrated pools almost always compensate LPs for loss-versus-holding, but trading fees are not sufficient to compensate for loss-versus-staking.
The analysis also includes two pools with reward-LSTs – rETH and cbETH – deployed to Curve v2. As reward-LSTs, rETH and cbETH are not pegged to 1 ETH and their pools use the Cryptoswap Invariant - an AMM that concentrates liquidity around their current prices. The returns in the rETH and cbETH pools were significantly impacted by the distribution of CRV token rewards. As depicted in Figure 8(a) (blue dotted lines), the LP wealth including CRV tokens actually significantly exceeded the staking returns for rETH and cbETH. Without rewards, the picture is similar to stETH pools, where trading fees are not sufficient to compensate for loss-versus-staking.
5.2.4 LST pools on Uniswap.
Figures 8(b), and 9(b) show the wealth of LP, HOLD, and LST, as well as the 30-day moving average of wealth difference, for the four LST pools on Uniswap v3. Due to incompatibility with rebase tokens, all LSTs listed at Uniswap are reward-based (wstETH, rETH, and cbETH). The moving averages plotted in Figure 9(b) shows that the pools with the higher fee rate of 5 basis points were able to generate a sufficient amount of fees to compensate for LVS during certain time periods, while the 1 basis point wstETH-ETH pool consistently failed to achieve this.
Over the whole observation period, total fee returns were generally high enough to compensate LPs for LVS in the 5 basis point pools, especially in the more recent past (see right plots in Figure 8(b)). However, the actual profitability of liquidity positions depends on their holding period, as it is affected by price fluctuations of cbETH and rETH. In particular, depositing into the pool during the depegs following the FTX insolvency in November 2022 (rETH being overpriced while cbETH was underpriced), can lead to different results.
@@ 5.3 Discussion
5.3.1 CRV rewards.
We find that most Curve pools currently are only profitable for LPs, especially compared to fully staking, when taking the CRV token rewards to LPs into account. The amounts of CRV token rewards emitted to pools are decided by the Curve governance, and can change every week. A drop in their distribution to the LST pools, could lead to lower returns for LPs and subsequently a decrease in liquidity in these pools.
5.3.2 CLMM Rebalancing.
@@
5.3.3 Reward-LST token pairs.
@@
6 Conclusion
Liquid staking tokens as liquidity on AMMs has grown to a vital part of the DeFi landscape, but has so far seen limited academic attention. This work systemizes liquidity provision of LSTs on AMMs, and both theoretically and empirically studies its returns.
In addition to LVH (“impermanent loss”), the most commonly used metrics to measure LP losses on AMM, we introduce loss-versus-staking (LVS) which is specifically tailored to the case of LST pools on AMMs. Our empirical results indicate that most AMM pools with LSTs have historically not compensated for loss-versus-staking. Nevertheless, most pools did compensate for LVH (“impermanent loss”).
This could lead to less LST liquidity being provided to AMMs in the long term. Currently, LST liquidity is already low compared to the overall LST supply (\$0.5 billion vs. over \$30 billion). A further decrease could negatively impact the price stability of LSTs and increase their risk of depegging.